OpenStudy (anonymous):

find the surface area of the part of the sphere x^2+y^2+z^2=9 that lies above the plane z=2 I found 21V21pi/6-pi/6, but I am not sure.

OpenStudy (perl):

so thats a sphere with radius 3 centered at origin

OpenStudy (anonymous):

yeah, is the answer right?

OpenStudy (aum):

Probably easiest to do this in spherical coordinates.

ganeshie8 (ganeshie8):

or polar coordinates maybe

OpenStudy (aum):



OpenStudy (aum):

phi goes from 0 to arccos(2/3) theta goes from 0 to 2pi

OpenStudy (aum):

\[\large \int_0^{\arccos(2/3)}\int_0^{2\pi}r^2\sin(\phi)d\phi d\theta \]

OpenStudy (anonymous):

thank you so much aum and ganeshie8. the hardest part for me was to write an equation. I got it. I can do the rest. thanks.

OpenStudy (aum):

I am getting \(\large 6\pi\) as the surface area.

ganeshie8 (ganeshie8):

same! same !

OpenStudy (anonymous):

by polar coordinates, i found different boundaries. dr, 0 to V3 and dØ, 0 to 2 pi.

OpenStudy (anonymous):

d delta, 0 to 2 pi

ganeshie8 (ganeshie8):

I see where you're doing the mistake

ganeshie8 (ganeshie8):

you need to look at the shadow in xy plane

ganeshie8 (ganeshie8):

you need to look at the shadow in xy plane of the surface u want

OpenStudy (anonymous):

is that x^2+y^2=5?

ganeshie8 (ganeshie8):

exactly !

ganeshie8 (ganeshie8):

it should be : x^2 + y^2 <= 5

ganeshie8 (ganeshie8):

its a disk

ganeshie8 (ganeshie8):

thats the region over which you need to evaluate teh double integral

OpenStudy (anonymous):

so, my answer was right? 21V21pi/6-pi/6

ganeshie8 (ganeshie8):

you should get 6pi i don't understand V's in your answer

OpenStudy (anonymous):

it is square root of 5 and I took the boundaries, dr, 0 to squared root of 5 and d(deta) 0 to 2pi. Am I doing something wrong?

OpenStudy (anonymous):

I found where I made the mistake. I found 6pi, too.

OpenStudy (anonymous):

thank you so much, ganeshie8. I appreciate you for helping me. Thank you.